OpenStudy (anonymous):

(-2)+(-3)(-4)-(3)(2)/(-2)(5)

OpenStudy (anonymous):

use the order of operations: (-2)+(12)-(-15) =25

OpenStudy (gw2011):

Evaluate the numbers within the parentheses first. This will give you the following: -2+12-(6)/(-10)=-2+12+6/10=106/10 (using the common denominator of 10). Changing this improper fraction to a decimal gives the answer of 10.6